A random sample of 146 recent donations at a certain blood bank reveals that 83 were type A blood. Does this suggest that theZ and P-Value needed. double-check answers thank you!
Answer #1

Solution :

Given that,

p_{0} = 0.40

1 - p_{0} = 0.60

n = 146

x = 83

Level of significance = \alpha = 0.01

Point estimate = sample proportion = \hat p = x / n = 0.569

This a two- tailed test.

The null and alternative hypothesis is,

Ho: p = 0.40

Ha: p \neq 0.40

Test statistics

z = (\hat p - p_{0} ) / \sqrt{} p_{0}*(1-p_{0}) / n

= ( 0.569 - 0.40) / \sqrt{} (0.40*0.60) / 146

= 4.16

P-value = 2 * P(Z > z )

= 2 * ( 1 - P(Z < 4.16))

= 2 * ( 1 - 1 )

= 0